OpenStudy (<3pandas18):

will give medal step by step plse. What is the solution of the linear-quadratic system of equations? {y = x^2 + 5x - 3} {y - x = 2}

OpenStudy (<3pandas18):

this is my guess x+2 = x^2+5x-3 x^2 + 4x - 5 = 0 (x+5)(x-1) = 0 x = -5 or 1 so, y = -3 or 3 =(-5,-3)(1,3)